OpenStudy (anonymous):

The distance traveled by a falling object is given by the formula d = 0.5gt2 where d = distance, g = the force of gravity, and t = time. Solve this equation for g, and use your formula to determine the force of gravity if a baseball takes 10 seconds to hit the ground after being dropped from a height of 480 feet. Show all steps in your work.

OpenStudy (mertsj):

1. Divide both sides of the equation by .5t^2 2. Replace d with 480. Replace t with 10. Simplify.

OpenStudy (anonymous):

don't you have to subtract ut from both sides

OpenStudy (mertsj):

Which part of "divide both sides by .5t^2" do you not understand?

OpenStudy (mertsj):

\[\frac{d}{.5t^2}=\frac{0.5t^2g}{.5t^2}\]

OpenStudy (mertsj):

\[\frac{d}{.5t^2}=g\]